If I have the ordered pairs of (2,54) &(54,2) How would I figure out where they would intersect. I know there is an equation for it but I can't remember hwo to get it started.
Answer by Alan3354(69443)

If I have the ordered pairs of (2,54) &(54,2) How would I figure out where they would intersect. I know there is an equation for it but I can't remember hwo to get it started.
------------------
First find the slope of the line thru the points.
slope m = diffy/diffx = -52/52 = -1
Then use either point and y = mx + b to find the y-intercept, b.
54 = -1*2 + b
b = 56
--> y = -x + 56
x+y = 56